# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 9.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.123123123Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 9.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.124124124If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 9.3. Even with hundreds of features, if only a handful matter, learning is fast. These results along with Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) establish the baseline rates for how structural assumptions affect learning speed in dynamic pricing.
@@ 9.2 Revealed Preference and Partial Identification
Misra et al. (2019) bring economic theory into the bandit pricing framework. Their model has $K$ discrete prices, $S$ consumer segments (segment membership observed, but valuations unknown), and within-segment heterogeneity $\delta$. A consumer in segment $s$ has valuation $v_{i}=v_{s}+n_{i}$, where $v_{s}$ is the segment midpoint and $n_{i}\in[-\delta,\delta]$ is idiosyncratic noise. The key structural assumption is WARP (Weak Axiom of Revealed Preference): if a consumer buys at price $p$, she would buy at any lower price $p^{\prime}<p$. This is equivalent to downward-sloping demand.
@@
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.125125125Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret:
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.126126126The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.127127127For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 9.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.128128128The regret benchmark here differs from Section 9.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).129129129EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).130130130The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. This separation quantifies, in regret units, the value of structural assumptions that econometricians routinely make. Knowing $F$ is the pricing analogue of specifying the structural model; not knowing $F$ is a more reduced-form agnostic position. When an economist specifies a logit or probit demand model, the assumed noise distribution is not merely a convenience for estimation; it purchases a qualitative improvement in the rate at which the algorithm learns. Conversely, semiparametric approaches that leave the error distribution unspecified pay a concrete cost in learning speed, reverting from logarithmic to polynomial regret.131131131Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.
@@ 9.4 Strategic Buyers
where $A$ is a positive definite matrix governing manipulation costs.132132132The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
@@
This is not $\sqrt{T}$ or $\log T$; it is linear in $T$. The revenue gap grows without bound and never stabilizes. Every standard dynamic pricing algorithm, including EMLP and RMLP, falls into this category because none accounts for strategic manipulation. The non-strategic policy is systematically biased. It underprices because it bases decisions on manipulated features, and the bias does not shrink with more data because the manipulation is endogenous to the pricing rule.
@@
Ignoring strategic behavior in online pricing is quite suboptimal. Incentive compatibility matters even in settings where the seller is “just” learning demand.133133133Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 9.5 Comparison of Regret Rates
Table 13 collects the regret rates across the papers discussed above. The table is ordered from weakest to strongest assumptions on demand, making visible three patterns. First, stronger assumptions yield faster learning, from $\sqrt{T}$ with no structure to $\log T$ with parametric or revealed-preference structure to $d\log T$ with known noise in high dimensions. Second, not knowing the noise distribution is expensive; moving from known to unknown $F$ costs a super-polynomial factor in the regret rate. Third, strategic behavior destroys standard algorithms, producing linear regret that no amount of data can overcome without an explicit correction.
@@ 9.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge. The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus. Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds.
Table 14 and Figure 6 report cumulative regret at four checkpoints. The results confirm the theoretical progression from Table 13. $\varepsilon$-greedy regret grows linearly ($R/T$ stabilizes near 0.011). LTE pays its exploration cost upfront, with regret of 1,005 at T=10, [^0] T=10{,}000 (entirely during the random phase), but flattens as the committed price earns near-optimal profit; $R/T^{2/3}$ stabilizes near 0.5. UCB1 and Thompson Sampling both grow as $\sqrt{T}$, with Thompson achieving a five-fold better constant (2.5 vs. 15.1 in $R/\sqrt{T}$ at T=200, [^0] T=200{,}000). UCB-PI-tuned achieves the lowest regret at every checkpoint past T=10, [^0] T=10{,}000, reaching 780 at T=200, [^0] T=200{,}000, a 31% reduction over Thompson Sampling and a 83% reduction over UCB-PI without variance tuning. The untuned UCB-PI variant overexplores; its regret of 4,503 scales as $\sqrt{T}$ rather than the $\log T$ that the theoretical bound predicts with correct tuning, placing it closer to UCB1 (6,734) than to the tuned variant (780).

### 15 Mar 2026 v2

@@ 11.1.3 High-Dimensional Features with Sparsity
Javanmard and Nazerzadeh (2019) extend the setting to products described by $d$-dimensional feature vectors $x_{t}$. The market value is $v_{t}=x_{t}^{\top}\theta_{0}+z_{t}$, where $\theta_{0}\in\mathbb{R}^{d}$ is unknown and $z_{t}$ is i.i.d. noise from a known log-concave distribution FF.174174174Log-concavity of $F$ and 1F1-F is satisfied by the normal, logistic, uniform, Laplace, and exponential distributions. It ensures that expected revenue $p\cdot[1-F(p-x_{t}^{\top}\theta_{0})]$ is strictly quasi-concave in $p$, giving a unique optimal price. Only $s_{0}$ of the $d$ coordinates of $\theta_{0}$ are nonzero, but the seller does not know which ones.
Their RMLP algorithm (Regularized Maximum Likelihood Pricing) operates in episodes whose lengths double (1, 2, 4, 8, …). At each episode boundary, the seller fits a LASSO-penalized maximum likelihood estimate of $\theta_{0}$ using data from the previous episode, then prices greedily throughout the current episode. The regret is $O(s_{0}\log d\cdot\log T)$ (Theorem 4.1), with a matching lower bound of $\Omega(s_{0}(\log d+\log T))$ (Theorem 5.1). The reason this beats $\sqrt{T}$ connects back to the Broder lower bound. In the parametric model of Section 11.1.2, all demand curves can cross at the optimal price, making that price uninformative. Here, customer features vary across periods, so the aggregate demand function at any fixed price changes in proportion to the estimation error in $\theta_{0}$. Every price is informative, and dedicated exploration rounds become unnecessary.175175175If some feature directions are rarely observed, the seller cannot learn all coordinates of $\theta_{0}$ quickly, and regret degrades to $O(\sqrt{\log(d)\cdot T})$ (Theorem 4.2). If the noise distribution belongs to a known parametric family but its scale parameter is unknown, regret reverts to $\Omega(\sqrt{T})$ (Theorem 7.1), foreshadowing the result of Xu and Wang (2021) discussed in Section 11.3. Even with hundreds of features, if only a handful matter, learning is fast.
@@ 11.2 Revealed Preference and Partial Identification
Misra et al. (2019) bring economic theory into the bandit pricing framework. Their model has $K$ discrete prices, $S$ consumer segments (segment membership observed, but valuations unknown), and within-segment heterogeneity $\delta$. A consumer in segment $s$ has valuation $v_{i}=v_{s}+n_{i}$, where $v_{s}$ is the segment midpoint and $n_{i}\in[-\delta,\delta]$ is idiosyncratic noise. The key structural assumption is WARP (Weak Axiom of Revealed Preference): if a consumer buys at price $p$, she would buy at any lower price $p^{\prime}<p$.
@@
if $p_{k}$ is not dominated, and $I_{p_{k}}(t)=0$ otherwise, where $\hat{\mu}_{p_{k}}(t)$ is the average profit observed at price $p_{k}$ and $N_{p_{k}}(t)$ is the number of trials.176176176Standard UCB1 uses an exploration bonus of $\sqrt{2\ln t/N_{p_{k}}(t)}$, which assumes rewards in $[0,1]$. Since profit at price $p_{k}$ is bounded by $p_{k}$, scaling the bonus by $p_{k}$ tightens exploration for cheap prices that cannot contribute much profit regardless. Two innovations drive the improvement over standard UCB1. First, dominance elimination reduces the effective number of arms the algorithm must explore. Second, the $p_{k}$ scaling focuses exploration on prices where uncertainty actually matters for profit. Together, these yield $O(\log T)$ regret.
@@
where $\Delta_{k}=\mu_{k^{*}}-\mu_{k}$ is the gap between arm $k$ and the optimal arm.177177177The first sum is the leading term, scaling as $O(\log T)$. The second sum is a constant that does not grow with $T$. Replacing $p_{k}$ with 1 recovers the standard UCB1 bound, which is looser.
Misra et al. (2019) calibrate the model to a field experiment at ZipRecruiter, a B2B online recruiting platform. With 7,870 customers per month, 1,000 segments, and 10 price points from \$19 to \$399, UCB-PI achieves 98% of oracle profit and produces 43% higher profits during the first month of testing compared to a learn-then-earn alternative. The algorithm has both higher mean profit and lower variance, reflecting the value of eliminating dominated prices early.178178178For multi-product settings, Mueller et al. (2019) impose low-rank structure on the price-sensitivity matrix, achieving regret $O(T^{3/4}\sqrt{d})$ that scales with the latent demand dimension $d$ rather than the number of products. Badanidiyuru et al. (2013) extend the bandit framework to handle inventory constraints (“bandits with knapsacks”), relevant when the seller faces limited stock alongside the pricing decision.
@@ 11.3 The Value of Knowing the Noise Distribution
Xu and Wang (2021) ask how much it helps to know the shape of demand uncertainty. In their model, a feature vector $x_{t}\in\mathbb{R}^{d}$ describes each sales session, the customer’s valuation is $w_{t}=x_{t}^{\top}\theta^{*}+N_{t}$ where $\theta^{*}$ is unknown and $N_{t}$ is zero-mean i.i.d. noise with CDF $F$, and the seller observes only whether the customer bought at the posted price. The answer depends on whether $F$ is known.179179179The regret benchmark here differs from Section 11.1.1. Kleinberg and Leighton (2003) and Broder and Rusmevichientong (2012) measure regret against the best fixed price; Xu and Wang (2021) and Liu et al. (2024) measure regret against the clairvoyant contextual policy that sets the optimal price $p_{t}^{*}$ for each customer’s features $x_{t}$. The contextual benchmark is harder, which makes the $O(d\log T)$ rate more impressive.
@@
If $F$ is known (the seller knows that demand shocks are, say, normally distributed with known variance), the EMLP algorithm (Epoch-based Maximum Likelihood Pricing) achieves regret $O(d\log T)$ (Theorem 3).180180180EMLP runs in doubling epochs of length $\tau_{k}=2^{k-1}$. At each epoch boundary, the seller fits a maximum likelihood estimate $\hat{\theta}_{k}$ using data from the previous epoch, then prices greedily at $p_{t}=J(x_{t}^{\top}\hat{\theta}_{k})$ throughout the epoch, where $J(u)=\arg\max_{v}\,v[1-F(v-u)]$ is the revenue-maximizing price function. The key technical insight is that the negative log-likelihood is strongly convex (Lemma 7 of Xu and Wang (2021)), so MLE concentrates at rate $O(d/\tau_{k})$. Since regret is quadratic in the parameter estimation error (Lemma 5) and there are $O(\log T)$ epochs, the total regret is $O(d\log T)$. After 10,000 customers with $d=5$ features, the revenue loss is roughly 50 customers’ worth, an improvement over the $\sqrt{T}\approx 100$ customers’ worth that Kleinberg’s lower bound imposes without structural knowledge.
@@
If $F$ is unknown (even if only the variance of a Gaussian is unknown, with everything else known), the regret is at least $\Omega(\sqrt{T})$ (Theorem 12).181181181The lower bound constructs two noise variances $\sigma_{1}=1$ and $\sigma_{2}=1-T^{-1/4}$. Any algorithm that performs well under both must spend $\Omega(\sqrt{T})$ revenue distinguishing the two cases. This extends the “uninformative price” phenomenon of Broder and Rusmevichientong (2012): when the seller does not know $F$, there exist prices at which observed purchase behavior is nearly identical under different demand parameters. The seller is back to the Kleinberg baseline, with no algorithm able to achieve sublinear improvement regardless of how many features are available.
@@
The gap between $O(d\log T)$ and $\Omega(\sqrt{T})$ is super-polynomial in $T$, not merely a constant factor. When an economist specifies a logit or probit demand model, the assumed noise distribution purchases a qualitative improvement in learning rate. Semiparametric approaches that leave the error distribution unspecified pay a concrete cost, reverting from logarithmic to polynomial regret.182182182Tullii et al. (2024) establish the tightest known bound under minimal distributional assumptions: if the noise distribution (c.d.f.) is merely Lipschitz continuous, the minimax regret is $\Theta(T^{2/3})$, strictly between the $\log T$ rate with known $F$ and the $\sqrt{T}$ rate with unknown $F$. Fan et al. (2024) consider a semiparametric setting where the noise density is smooth and connect the pricing problem to the econometrics of semiparametric estimation.
@@ 11.4 Strategic Buyers
where $A$ is a positive definite matrix governing manipulation costs.183183183The matrix $A$ captures how costly it is for the buyer to distort each feature dimension. High eigenvalues of $A$ mean manipulation is expensive. This is the standard model of strategic classification (Hardt et al., 2016), adapted to pricing. The first-order condition yields a systematic bias: the buyer shifts her features to make the seller’s model predict a lower valuation, securing a lower price. The seller observes only the distorted features $\tilde{x}_{t}$, not the true $x_{t}^{0}$.
@@
The regret is linear in $T$: every standard dynamic pricing algorithm, including EMLP and RMLP, systematically underprices because it bases decisions on manipulated features, and the bias does not shrink with more data because the manipulation is endogenous to the pricing rule.
@@
Incentive compatibility matters even in settings where the seller is “just” learning demand.184184184Agrawal and Tang (2024) document a related phenomenon in pricing with reference effects. If consumers anchor on past prices, a static pricing policy that ignores reference dependence accumulates linear regret $\Omega(T)$. Chen et al. (2025) show that imposing fairness constraints (requiring similar prices for similar customers) raises the regret floor to $\Theta(T^{2/3})$, a social cost of equitable treatment.
@@ 11.5 Comparison of Regret Rates
Table 21 collects regret rates ordered from weakest to strongest structural assumptions. The dominant pattern is that stronger assumptions yield faster learning, with the gap between $\log T$ and $\sqrt{T}$ being super-polynomial in $T$, not merely a constant factor. Strategic behavior is the outlier: it produces linear regret that no amount of data can overcome without explicit correction. Figure 19 plots each rate on a log-log scale.
@@
Figure 19: Theoretical regret rate functions at constants equal to 1, $d=5$. Two cases for the $s_{0}\log d\log T$ rate are shown: $s_{0}=1$ (very sparse, $\approx 15$ lost at T=10, [^0] T=10{,}000) and $s_{0}=5$ (moderate sparsity, $\approx 74$ lost). The vertical line marks T=10, [^0] T=10{,}000, matching the “Per 10K” column of Table 21.
@@ 11.7 Simulation Study: The Knowledge Ladder
I run six algorithms on the Misra et al. (2019) demand environment to trace how cumulative regret responds to increasing structural knowledge.185185185The environment has $K=100$ prices on a grid from \$0.01 to \$1.00, S=1, [^0] S=1{,}000 consumer segments with equal weights, within-segment heterogeneity $\delta=0.1$, and segment midpoints $v_{s}\sim\mathrm{Uniform}(0.1,0.9)$. A consumer purchases if and only if $v_{i}\geq p$ (WARP). Each algorithm runs across 10 seeds with T=200, [^0] T=200{,}000 rounds. The six algorithms, ordered by the structural knowledge they exploit, are: (0) $\varepsilon$-greedy with $\varepsilon=0.1$, which never adapts its exploration rate; (1) Learn-Then-Earn (LTE), which explores uniformly for the first 5% of rounds and then commits to the empirical best price; (2) UCB1 (Auer et al., 2002a), which adapts exploration via confidence bounds but ignores demand structure; (3) Thompson Sampling (Thompson, 1933), which maintains a Bayesian posterior over purchase rates186186186At each period the algorithm draws one sample $\tilde{\mu}_{k}$ from each arm’s posterior over its purchase rate and selects the arm with the highest sampled expected profit $p_{k}\tilde{\mu}_{k}$; arms with uncertain posteriors have high-variance draws and are selected frequently, while well-estimated arms are selected in proportion to how likely they are optimal.; (4) UCB-PI (Misra et al., 2019), which uses WARP to eliminate dominated prices and scales the exploration bonus by the price level; and (5) UCB-PI-tuned, which adds a variance-based refinement to the exploration bonus.
Table 22 and Figure 20 report cumulative regret at four checkpoints. The results confirm the theoretical progression from Table 21: $\varepsilon$-greedy grows linearly, UCB1 and Thompson Sampling grow as $\sqrt{T}$, and UCB-PI-tuned achieves the lowest regret at every checkpoint past T=10, [^0] T=10{,}000. The untuned UCB-PI variant overexplores, scaling as $\sqrt{T}$ rather than the theoretical $\log T$, illustrating that WARP-based elimination alone is insufficient without variance-calibrated exploration bonuses.